OpenStudy (anonymous):

A kilobyte is 2^10. A megabyte is 2^20 bytes. How many kilobytes are in a megabyte?

OpenStudy (anonymous):

divide the 2

OpenStudy (primeralph):



OpenStudy (anonymous):

1024 is 2^10 1048576 is 2^20 So i figured id do something with division... But idk what

OpenStudy (mathstudent55):

\(a^m \div a^n = a^{m - n} \) \(2^{20} \div 2^{10} = 2^{20 - 10} = 2^{10}\)

OpenStudy (anonymous):

Thank you all

OpenStudy (primeralph):

You're welcome.

OpenStudy (mathstudent55):

wlcm